Prepare a frequency distribution from the following data relating to bonus paid to factory workers.


86, 62, 58, 73, 101, 90, 94, 90, 76, 61, 84, 63, 56, 72, 102, 56, 83, 92, 87, 60, 83, 69, 57, 71, 103, 57, 87, 93, 88, 59, 76, 70, 54, 70, 104, 58, 88, 94, 89, 57, 74, 86, 55, 60, 105, 59, 89, 84, 90, 74, 67, 84, 82, 70, 60, 60, 90, 81, 91, 76, 60, 83, 78, 80, 65, 61, 96, 82, 93, 102, 60, 91, 76, 90, 70, 63, 94, 83, 94, 101, 70, 100, 74, 101, 80, 67, 92, 85, 96, 92, 60, 83, 78, 80, 65, 61, 96, 82, 93, 76.


Take a class of interval of five

1
Expert's answer
2022-01-10T11:39:44-0500

Class of interval (k) = 5

"No.of.Classes=\frac{Maximum-Minimum}{k}=\frac{105-54}{5}=10.2\approx11"

Following table shows a frequency distribution from the following data relating to bonus paid to factory workers: